CERN BT Report, Dec. 3rd 2012

Position scan with CREAM-2 HPD boxes in configuration 1 (front beam without lead) was completed.
Position scan with CREAM-2 HPD boxes in configuration 2 was processed until 8 am on Dec 3.
At 8 am on Dec 3 (Monday), beam stopped.
The instrument on the rotation fixture was moved from beam line to testing area.
Targets were unmounted from the pallet and stand-off under the target legs were removed.
TCD was removed. Targets were re-mounted with shims.
SBT and rotation fixture controller were uninstalled.
Client PC and (ethernet and serial) cables were uninstalled.
Packing started.

JiHye worked on removing LPSC's HPDs from ISS-CREAM HPD boxes
in order to disassemble an electronics board for HPD mounting on each HPD box.
Those boards were requested to be packed to UMD by MooHyun.
With the help of David, Mike, HyunGue and Kenny, wires on HPDs were cut with longest length possible, 
HPDs were un-mounted from HPD boxes, and the electronics boards were dis-assembled from the HPD box.
Electronics boards were packed into the crate.
LPSC's HPDs were put into a storage box one by one carefully.
Ludo picked up HPDs and HPD boxes to take LPSC on Dec. 4, Tuesday.
(When JiHye was trying to start working on HPD boxes, Jason pointed out that
she should follow the ESD training contents.
Since there was not ESD work bench and we did not bring ESD suit nor wrist strip,
we have to proceed it carefully. At least the table was grounded on the wall.)

Sasha helped with the KNU shipping arrangement to Korea.

 

-YS